A confined amount of gas at 10°C has pressure of 2.5atm. what is its new pressure if the temperature is increased to 50°C?
1
Expert's answer
2020-12-15T07:54:15-0500

"T_1 = 10°C = 273K + 10 = 283K\\
P_1 = 2.5atm"


"T_2 = 50°C = 273K + 50 = 323K\\
P_2 = ?"


Using Gay-Lussac's law,

P1/T1 = P2/T2

2.5/283 = P2/323

P2 = 2.5×323/283 = 2.853 atm


Therefore, the new pressure is 2.853 atm.